Solve using the substitution method. Show your work. If the system has no solution or an infinite number of solutions, state this.
-x + 6y = 54
3x + y = 9

Answer by mananth(16949)

-1 x + 6 y = 54 .............1
1 y = 9 -3 x ---------2
/ 1 =
y = 9 -3 x
Substitute y in (1)
-1 x + 6 ( 9 -3 x ) = 54
-1 x + 54 + -18 x = 54
-1 x -18 x = 54 -54
-19 x = 0
/ -19
x= 0
Plug the value of x in (1)
-1 x + 6 y = 54
-1 * 0 + 6 y = 54
0 + 6 y = 54
6 y = 54
/ 6
y= 9
Solution (0,9)
m.ananth@hotmail.ca